OpenStudy (anonymous):

Thirty Times The Square Of A Non-Zero Number Is Equal To 8 Times The Number What Is The Number

OpenStudy (anonymous):

30x^2 = 8x

OpenStudy (anonymous):

So, let x=the number we want to find, then we have: 30 * x^2 = 8 * x Divide both sides by x 30 * x = 8 Divide both sides by 30 and we have; x = 8/30 x = 4/15 x = 0.2666666......

OpenStudy (anonymous):

So The Number Is 0.2666666...........

OpenStudy (anonymous):

do u have any choices ?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

then yea dats what it is, but really it's 0.26